Most Liked Content
Posted by Guest on 04 October 2011 - 08:12 AM
domain name is under one reseller chain;
AND is using hosting from another reseller and is using that other resellers name servers;
Our admin team is already in the process of fixing this issue and we will promptly update you once it is fixed. Requesting your patience till then.
Posted by Guest on 14 January 2011 - 07:10 AM
While we have been working relentlessly to get the server back up online we have some bad news, the FSCK on the server has restarted. We predict that this is going to take another 10-12 hours, though we are not sure of the exact time. Further analysis confirms that it might be just that FSCK fixed errors during pass, but still found inconsistencies & hence it decided to check again.
We assure that we'll keep you posted about the update on a periodic basis so that you are not left in the dark.
Posted by Guest on 30 March 2012 - 04:21 PM
Start Time: 16.30:00 GMT
Users will experience service disruptions for imap/pop intermittently during this maintenance. Also new mails will be delayed.
For any additional information please contact our support desk.
Posted by Guest on 22 June 2011 - 02:27 PM
Since this fix is dependent upon the amount of data inconsistency, the time to apply the fix on various servers may vary.
Once the fix has propagated to all the servers, the email accounts should start to work. We will also keep you updated regarding the same in this thread.
Meanwhile, your patience is highly appreciated. For any further queries, please feel free to contact our support help-desk.
Posted by prashant.g on 02 February 2012 - 12:07 PM
This area is currently under Maintenance.
Please check again later.
The issue has been escalated and efforts are on to have this resolved as soon as possible.
Requesting your patience and co-operation in the meanwhile.
Posted by Guest on 15 March 2011 - 08:18 AM
The rsync of data from the old server to an alternate drive attached to the box overnight has completed successfully, however there are still some issue with the detection of the drive.
Below are the detailed steps that we did at our end to solve the issue.
* We rsycned all the data from the old server to an alternate drive attached to the box overnight
* This drive was moved to new server today but wasn't getting detected as on the old server it was attached to Adaptec RAID card and on the new server its connected to onboard.
* We have decided to start services on the old server and deal with frequent reboots till we migrate data over the network.
* But old lhb5 now seems to be suffering from frequent OOMs - we are attempting to get this fixed so that we can get the services back online.
Posted by Guest on 23 June 2011 - 08:15 AM
Hence we would be resetting the DB replication inorder to fix this.
Please find the details of the mentioned activity (DB replication) below:
Start time: 0900 hours GMT
End time: 1000 hours GMT
1) Mailbox Control Panel (old email architecture)
2) Login (including IMAP/POP)
Kindly note that there would be no mail loss. All the mails would be queued and will be delivered once the service is resumed.
Please watch this space for further updates.
We deeply regret the inconvenience caused due to this.
Posted by Guest on 06 May 2011 - 06:52 AM
We would keep the post updated once we have more details on what is the exact cause and the steps taken to avoid it.
Posted by priyanka.f on 25 March 2013 - 01:53 PM
Any requests hitting these six IP's may face an issue. However, the rest of our DNS IP's are working fine. We are closely monitoring to resolve this issue at the earliest.
We shall keep you updated on the same.
- Olunlade likes this
Posted by Guest on 12 September 2012 - 11:17 AM
We would keep you updated about the same.
In case of any further queries, feel free to contact the support helpdesk.
Inconvenience is highly regretted.
- dakine likes this
Posted by Guest on 15 January 2011 - 10:47 AM
The temporary workaround has been implemented and all the sites are resolving seamlessly. we still require to attach another standalone drive to the server and we are expecting this to be done by Monday.
In case there are any queries please feel free to contact our support desk.
- pablolm likes this
Posted by Guest on 28 December 2011 - 01:57 PM
Our Server Administration team is in the process of reconfiguring the sites so they can be up and accessible as soon as possible
477 sites from a total of 1046 have been reconfigured
Ideally all the sites should be up in a time span of 4-5 hours
This would depend on Plesk vhost reconfigurator which is currently running on the server
We appreciate your patience and cooperation in the meanwhile
Feel free to contact the support team for more details.
- Star Tech likes this
Posted by Guest on 24 November 2011 - 12:32 PM
A downtime has been scheduled for our Support Chat and Ticketing Interface for routine maintenance.
Date: Thursday November 24, 2011
Start Time : 0730 GMT
End Time : 0930 GMT
Chat Support will not be available throughout this time.You will also not be able to view your tickets.
However you can mail in your support requests and they will be queued up and replied to after the maintenance.
Phone support would be available through out this time span.
- firstname.lastname@example.org likes this
Posted by Guest on 24 September 2011 - 04:55 PM
The System Administration Team is working towards fixing this issue at the earliest.
We apologize for the inconvenience caused to you due to this issue.
Please watch this post for more updates.
- abinfocom likes this
Posted by Guest on 25 May 2011 - 06:45 AM
- Connectivity issues with the hosting control panels (via Orderbox)
- Slowness with Email services
- Slow resolution for the Websites
We deeply regret the inconvenience caused due to this event.
- Praveen likes this
Posted by Guest on 06 May 2011 - 02:19 PM
* We were hit by 1,20,000 qps per server
* Which was using 500Mbps bandwidth
Nearly 100/24 subnets (254000 ips) from china participated in that attack. All of these IP Addresses have been Blocked.
Please contact or Support Team if there you are still facing any particular issue.
- whdcentral likes this
Posted by Guest on 28 January 2015 - 07:17 AM
The problem with the server has now been identified and our system administration team is working to address the same at the earliest. An approximate ETA for the issue is 1 hour.
Please feel free to contact our support team in case you have any queries.
- WHYNOPE likes this
Posted by Guest on 23 January 2011 - 11:06 AM
Our System Administration Team has been informed abut this problem and they are working on fixing the same at the earliest.
Websites Hosted on the below mentioned IP Addresses seem to be affected.
We are deeply sorry for the inconvenience caused.
- Abdul likes this
Posted by Guest on 26 May 2011 - 08:02 AM
We identified the cause to be latency that was induced (under high traffic) by Juniper j4350 Routers/HP 2910Al aggregation switches that was being used at the Data Center. We have now replaced them with a Juniper Ex4200 series virtual chassis.
In case there are any further doubts, we would request you to kindly contact the support help desk at the earliest.
- Netcoderz likes this